OpenStudy (anonymous):

Which binomial is a factor of h^2 – 36 a. (h – 6) b. (h – 36) c. (h + 12) d. (h+ 36)

rebeccaxhawaii (rebeccaxhawaii):

any ideas?

OpenStudy (anonymous):

@rebeccaxhawaii h + 36?

zepdrix (zepdrix):

36 is a perfect square, and can be rewritten this way,\[\large\rm h^2-36\quad=\quad h^2-6^2\]Our `difference of squares` rules tells us how to factor,\[\large\rm a^2-b^2=(a-b)(a+b)\]